### v1

# Less Is More for Multi-Step Logical Reasoning of LLM Generalisation Under Rule Removal, Paraphrasing, and Compression

Abstract
Large language models (LLMs) excel across many natural language tasks, yet their generalisation to structural perturbations in logical contexts remains poorly understood. We introduce a controlled evaluation framework that probes reasoning reliability through four targeted stress tests: (1) rule deletion, removing either redundant or essential rules from a multi-step inference chain; (2) contradictory evidence injection; (3) logic-preserving rewrites generated through several families of equivalence laws (contrapositive, double negation, implication, De Morgan, identity, and commutativity); and (4) multi-law equivalence stacking that introduces 2–5 simultaneous logical transformations.
Across three representative model families—BERT, Qwen2, and LLaMA-like models—our experiments reveal a strikingly consistent pattern: all models achieve perfect accuracy on the base tasks and remain fully generalise to redundant rule deletion and all equivalence-based rewrites (single or multi-law), but fail sharply under essential rule deletion (dropping to 25% accuracy) and collapse completely in the presence of explicit contradictions (0% accuracy). These results demonstrate that LLMs possess stable invariance to semantic-preserving logical transformations, yet remain fundamentally brittle to missing or conflicting evidence. Our framework provides a clean diagnostic tool for isolating such reasoning failure modes and highlights persistent gaps in the logical generalisation abilities of current LLMs.
Codehttps://github.com/14H034160212/lemo
Introduction
Reasoning over structured rule systems is a fundamental capability required for reliable decision-making and language understanding. Although large language models (LLMs) achieve impressive performance across numerous natural language benchmarks, it is still unclear to what extent these models perform true logical reasoning rather than relying on memorized patterns or superficial correlations. Much prior work evaluates LLMs in settings where all relevant information is present in canonical form; however, real-world reasoning often involves incomplete information, rephrased or transformed rules, redundant evidence, and even explicit contradictions. Understanding how LLMs behave under such perturbations is essential for both theoretical insight and safe deployment.
To address this gap, we introduce a highly controlled evaluation framework that isolates core aspects of rule-based reasoning. The framework is built around a canonical multi-step logical chain (e.g., green → cold → rough → young → nice) that requires models to integrate several rules in sequence. From this base structure, we generate a family of targeted variants that probe different dimensions of reasoning generalisation:
Rule deletion. We remove either redundant rules—whose absence should not change conclusions—or essential rules that constitute the backbone of the inference chain. This tests whether models distinguish between logically important and unimportant components.
Contradictory evidence injection. We introduce explicit facts that conflict with conclusions implied by the rule chain, requiring the model to reconcile or prioritize between mutually inconsistent pieces of information.
Logic-preserving transformations. We rewrite rules using several major families of logical equivalence laws, including contraposition, double negation, implication-to-disjunction, De Morgan’s laws, identity, and commutativity of disjunction. These variants preserve the underlying semantics while changing surface form, allowing us to test whether LLMs track logical content rather than specific linguistic templates.
Multi-law equivalence stacking. We introduce conditions in which 2–5 logical equivalence laws are applied simultaneously to create highly redundant, cluttered, and structurally complex rule sets, all of which remain logically equivalent to the base rules.
We evaluate this framework across three distinct model families—BERT, Qwen2, and LLaMA-like models—using identical training and testing pipelines. Despite architectural differences, all models exhibit the same qualitative and quantitative behavior:
  • Perfect reasoning on the base tasks (100% accuracy)
  • Full generalisation to redundant rule deletion (100% accuracy)
  • Catastrophic failure under essential rule deletion (25% accuracy)
  • Complete collapse under explicit contradictions (0% accuracy)
  • Full invariance to all six classes of logical equivalence rewrites (100% accuracy)
  • Full invariance even under multi-law equivalence stacking (100% accuracy)
Together, these results reveal a consistent pattern: LLMs correctly identify which rules are structurally necessary, generalise flawlessly across a wide spectrum of logically equivalent transformations, but lack the capacity to adjust reasoning when evidence is incomplete or contradictory. Rather than flexibly integrating conflicting information, models tend to complete the rule chain whenever possible, even when doing so contradicts stated facts.
Our contributions provide a clean and extensible diagnostic tool for probing reasoning behavior, enabling precise identification of reasoning failure modes. The findings highlight a persistent gap between surface-level generalisation to paraphrase-like variation and broader generalisation to structural perturbation or conflict, underscoring the need for models and training strategies that can reason reliably under minimal, missing, or inconsistent evidence.
Related Work
Research on logical reasoning in large language models (LLMs) spans multi-step deduction, logic-driven augmentation, abductive explanation, abstract reasoning assessment, generalisation analysis, and hybrid neuro-symbolic methods. A growing body of work shows that LLMs often fail even on elementary logical transformations. For example, the Reversal Curse demonstrates that models trained on “A is B” frequently fail to infer the logically equivalent “B is A,” revealing structural weaknesses in bidirectional generalisation (berglund2023reversal). Such findings motivate deeper investigation into how LLMs internalise logical relationships and under which conditions reasoning fidelity deteriorates.
Early studies like Transformers as Soft Reasoners demonstrated that transformers can perform multi-step inference over natural language rules and generalise to deeper reasoning depths (transformerClark). Subsequent work showed that this generalisation is fragile: Bao et al. found that models trained on short reasoning chains fail to extrapolate to longer or structurally shifted chains, highlighting persistent deficiencies in compositional generalisation (bao2022multi). Similar concerns are echoed in broader analyses showing that LLMs lack strong inductive biases and struggle with systematic, relational, or abstract reasoning (ijcai2024p693; Cheng2025Em; cheng2026sapp).
A parallel line of work attempts to strengthen reasoning via logic-driven augmentation. Wang et al. propose a symbolic context-extension and augmentation framework that applies equivalence laws to derive new training instances, improving model performance on ReClor and LogiQA (wang-etal-2022-logic). Bao et al. extend this idea by using Abstract Meaning Representation (AMR) to generate semantically grounded logical variants that enhance generalisation to unseen reasoning depths (bao2024). Despite their effectiveness, these methods primarily expand data distributions rather than isolating the minimal perturbations that cause reasoning failures.
Beyond deduction, abductive reasoning has been explored through frameworks such as AbductionRules, which train models to generate rules explaining surprising inputs (young-etal-2022-abductionrules). Other studies examine LLM generalisation under structured perturbations: Bao et al. show that reordering, compressing, or restructuring premises leads to substantial degradation in reasoning fidelity (baoAccess). Hybrid neuro-symbolic approaches, such as ChatLogic, integrate LLMs with Prolog-style inference mechanisms to address these limitations, although they rely heavily on external symbolic scaffolds (10650138). Complementing these evaluations, Bao’s dissertation provides a unified perspective on out-of-distribution generalisation, structural generalisation, and equivalence invariance by analysing architecture-agnostic failure modes across premise deletion, contradiction insertion, and logic-preserving rewrites (bao2025thesis).
Across these threads, prior work demonstrates meaningful progress but still lacks frameworks capable of systematically disentangling generalisation to redundant versus essential evidence, contradiction sensitivity, and invariance to different classes of logical equivalence. These unresolved limitations motivate our controlled, fine-grained evaluation framework that isolates minimal structural perturbations and reveals predictable failure modes in multi-step reasoning across LLM architectures.
Methodology
1. Synthetic Logical Environment
Base facts + rule chain generation
$\Downarrow$
2. Structural Variants
V1: Redundant rule removal
V2: Essential rule removal V3: Contradictory facts V4: Logical equivalence laws (1–5)
$\Downarrow$
3. Model Fine-Tuning & Inference
BERT / Qwen2 / TinyLlama (LoRA)
$\Downarrow$
4. Evaluation & Analysis
Accuracy, $\Delta_{\text{base}}$, variant breakdown
Figure 1: Compact system architecture of our logical reasoning generalisation framework. Each base example is transformed into structural variations (rule deletion, contradiction injection, and logic-equivalence rewrites). Models are trained on base data only and evaluated across all variants to measure generalisation and sensitivity.
The system architecture in Figure 1 outlines the full pipeline used to construct, transform, and evaluate logical reasoning tasks. The process begins with the generation of synthetic logical environments containing base facts and a multi-step rule chain designed to support deterministic deductive inferences. From each base environment, we systematically derive multiple structural variants that probe different dimensions of reasoning generalisation: (1) removal of redundant rules, (2) removal of essential rules, (3) injection of explicit contradictions, and (4) application of a wide range of logical-equivalence transformations.
These transformed instances, together with the untouched base examples, form a comprehensive evaluation suite. Models such as BERT, Qwen2, and TinyLlama are fine-tuned solely on the base cases using LoRA-based parameter-efficient adaptation. They are then tested on all structural variants to assess reasoning generalisation under controlled perturbations. The evaluation module computes accuracy and deviation from the base condition ($\Delta_{\text{base}}$), enabling fine-grained analysis of model generalisation, sensitivity to missing or conflicting information, and stability under logic-preserving rewrites.
Extended Logic Rules for Multi-Step Reasoning
This chapter explores the reasoning capabilities of AI through a comprehensive case study that demonstrates the principle of “less is more” in logical deduction. We begin with a complex base example incorporating dilemma reasoning patterns, then systematically modify it in three distinct ways to reveal different aspects of AI reasoning:
  1. 1.
    Rule Reduction: Removing redundant rules while maintaining the same conclusions
  2. 2.
    Rule Equivalence: Replace multiple rules with fewer logically equivalent rules, including the laws of contraposition, implication, double negation, De Morgan, identity, and commutativity.
  3. 3.
    Rule Interference: Adding contradictory rules that disrupt the AI’s reasoning process.
All examples use consistent predicates and individuals (Anne, green, blue, cold, rough, young) to facilitate comparison across variations. The example for each variation can be found in the appendix.
Experiment
Dataset Construction To evaluate reasoning generalisation under controlled structural perturbations, we construct a synthetic dataset consisting of multi-step logical inference chains and systematically generated variants. Each base example includes a fact specifying an entity’s attribute (e.g., “Anne is green or blue”), a rule chain of five inference steps (e.g., green → cold → rough → young → nice), four questions querying intermediate and final states, and binary ground-truth answers. From each base example, we generate eleven diagnostic variants grouped into four categories:
Rule Deletion. Variant 1: removal of a logically redundant rule. Variant 2: removal of a logically essential rule. Contradictory Evidence. Variant 3: addition of an explicit contradiction in the facts. Logic-Preserving Equivalence Laws. Six single-law variants using distinct logical equivalence transformations: contrapositive, double negation, implication, De Morgan, identity, and commutativity. Multi-Law Equivalence Stacking. A final variant consisting of 2–5 equivalence-law transformations simultaneously applied to the base rule.
The dataset contains 100 base groups, of which 80 are used for training and 20 for testing; all variants are applied to all base groups, resulting in a comprehensive evaluation suite with consistent ground truth. All transformations retain natural-language surface forms to ensure that evaluation probes reasoning rather than symbolic manipulation skills.
Models Evaluated We evaluate three representative model families commonly used in reasoning research: BERT-base (devlin2018bert): a bidirectional encoder-only architecture. Qwen2-1.5B (yang2025qwen3): a decoder-only transformer representative of modern instruction-tuned LLMs. TinyLlama-1.1B (zhang2024tinyllama): an efficient LLaMA-like decoder model trained on open datasets.
Training Procedure All models are fine-tuned using LoRA (Low-Rank Adaptation) to ensure parameter-efficient training and allow direct comparison across architectures. Training is conducted on the 80 base examples only—none of the variants are seen during training. Each question is treated as an independent binary classification instance.
Input Formatting Each training instance concatenates: [facts] + [rules] + [question] with the binary label corresponding to the ground-truth answer (“T” or “F”).
Hyperparameters To maintain comparability across models, all fine-tuning uses: Batch size: 4, Learning rate: $2\times 10^{-5}$, Epochs: 10, AdamW optimizer, Maximum sequence length: 512, and LoRA rank: 8. For decoder-only models (Qwen2 and TinyLlama), the tokenizer’s pad_token is set to the eos_token to avoid padding-related errors.
Evaluation Protocol We evaluate each model on all eleven test splits: Base test set (20 unseen base examples), Rule Deletion Variants, Contradictory Evidence Variant, Six Equivalence-Law Variants, and Multi-Law Variant.
Each test contains four questions per example, producing 80 questions for the base set and 400 questions per variant. For each input, the model predicts a binary label.
During evaluation, we additionally record: The exact facts and rules presented, the equivalence laws used (if any), Question-level predictions, and a human-readable description of the structural perturbation. This allows granular error analysis of different reasoning stressors.
We use accuracy as the primary metric, reflecting the percentage of correctly answered questions in each test split. Because each question is binary and independent, accuracy provides a direct measure of reasoning correctness.
Discussion and Implications
BERT Qwen2 TinyLlama
Split Acc $\Delta$ Acc $\Delta$ Acc $\Delta$
base 1.00 0.00 1.00 0.00 1.00 0.00
variant1 1.00 0.00 1.00 0.00 1.00 0.00
variant2 0.25 -0.75 0.25 -0.75 0.25 -0.75
variant3 0.00 -1.00 0.00 -1.00 0.00 -1.00
variant4-contrapositive 1.00 0.00 1.00 0.00 1.00 0.00
variant4-double-negation 1.00 0.00 1.00 0.00 1.00 0.00
variant4-implication 1.00 0.00 1.00 0.00 1.00 0.00
variant4-de-morgan 1.00 0.00 1.00 0.00 1.00 0.00
variant4-identity 1.00 0.00 1.00 0.00 1.00 0.00
variant4-commutativity 1.00 0.00 1.00 0.00 1.00 0.00
variant4-multi (2–5 laws) 1.00 0.00 1.00 0.00 1.00 0.00
Table 1: Accuracy and deviation from base ($\Delta$) for all models across structural variants.
Table 1 presents a unified comparison of BERT, Qwen2, and TinyLlama across all structural variants in our benchmark. Despite architectural differences, all three models exhibit an almost perfectly aligned pattern of behavior. Each model achieves 100% accuracy on the base reasoning task and retains this performance under redundant rule removal (Variant 1) as well as all logic-preserving equivalence transformations, including six single-law rewrites and multi-law (2–5) equivalence stacking. These conditions yield $\Delta=0$, confirming complete invariance to semantic-preserving logical reformulations.
In contrast, performance collapses when essential rule information is removed (Variant 2), with accuracy dropping uniformly to 25% for all models, corresponding to a $\Delta$ of 0.75-0.75. Even more strikingly, all models fail entirely under contradictory facts (Variant 3), scoring 0% accuracy and yielding the largest degradation from baseline ($\Delta=-1.00$). This cross-model consistency indicates that generalisation to logical equivalence and brittleness to structural disruptions are not tied to model size or architecture, but instead reflect systematic properties of current LLM reasoning mechanisms. The variations in our framework reveal key properties of LLM-based reasoning. By systematically manipulating the logical environment—removing rules, introducing contradictions, and applying logic-preserving equivalence transformations—we expose dimensions of generalisation and brittleness often hidden in more complex benchmarks. Together, these findings clarify what contemporary LLMs can and cannot generalise in rule-based reasoning.
Generalisation and Sensitivity
The progression from identical conclusions (Variation 1) to partially disrupted conclusions (Variation 2) to complete breakdown (Variation 3) reveals the fine-grained sensitivity of LLM reasoning to structural properties of the rule set. In Variation 1, the removal of a redundant rule leaves all models completely unaffected—indicating that LLMs can identify when a rule does not contribute new inferential content. In sharp contrast, Variation 2 demonstrates that eliminating a single essential step in the reasoning chain causes accuracy to collapse to chance levels, suggesting that LLMs cannot reconstruct missing inferential links even when the rest of the context is intact.
Variation 3 highlights a different form of fragility: all models overwhelmingly prefer rule-driven inferences over explicitly stated contradictory facts. This behavior implies that LLMs tend to follow the internal logic of the given rule system rather than adjusting their conclusions in response to new or conflicting information—an important consideration for applications involving dynamic or noisy environments.
The invariance observed across all logic-preserving rewrites—whether single-law or multi-law—shows that LLMs possess semantic stability, treating contraposition, double negation, implication re-expression, De Morgan transformations, and other equivalence laws as interchangeable. This indicates that they capture structural relationships between logically equivalent formulations, even though they falter with incomplete or inconsistent evidence. Overall, LLMs remain generalise to semantic-preserving transformations but extremely brittle to structural disruptions, a duality with important implications for AI safety: systems that appear reliable under surface-level variation may still fail when essential information is missing or contradictory, highlighting the need for methods that compensate for incomplete reasoning or reconcile conflicting evidence.
Conclusion
This work introduced a controlled framework for evaluating LLM generalisation under systematic perturbations to logical reasoning. By constructing multi-step inference chains and generating variants involving rule deletion, contradiction injection, and multiple families of logic-preserving transformations, we isolated reasoning behaviors often obscured in prior benchmarks. Across BERT, Qwen2, and LLaMA-like models, we observed perfect performance and complete invariance to semantic-preserving rewrites, but sharp failures when essential rules were removed and total collapse under explicit contradictions. These results highlight a clear gap between surface-level generalisation and deeper logical competence, underscoring the need for models that can integrate incomplete information, handle conflicting evidence, and recover missing reasoning steps. Our framework provides a foundation for improving both LLM reasoning reliability and evaluation methodology.
Appendix A Appendix
Base Example: Complex Dilemma Reasoning Structure
We begin with a comprehensive example that establishes the core reasoning pattern, drawing on a structure analogous to the Paradox of the Court [^2] The Paradox of the Court involves a contract between the teacher Protagoras and his student Euathlus, where the student only pays for lessons if he wins a court case. When Protagoras sues Euathlus for the fee, a paradox arises: if Euathlus wins, he owes nothing, but if he loses, he still avoids payment, creating a logical contradiction about the outcome., a classic logical dilemma where multiple possible paths lead to the same conclusion, despite their apparent differences.
  • Facts:
    • Anne is green or blue
  • Rules:
    • Rule 1: If someone is green then they are cold. $\forall x\,(\text{Green}(x)\rightarrow\text{Cold}(x))$
    • Rule 2: If someone is blue then they are cold. $\forall x\,(\text{Blue}(x)\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is not young then they are not rough. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
    • Rule 5: If someone is young then they are cold. $\forall x\,(\text{Young}(x)\rightarrow\text{Cold}(x))$
    • Rule 6: If someone is young then they are nice. $\forall x\,(\text{Young}(x)\rightarrow\text{Nice}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: T]
    • Q2: Anne is rough. True/False? [Answer: T]
    • Q3: Anne is young. True/False? [Answer: T]
    • Q4: Anne is nice. True/False? [Answer: T]
The fact “Anne is green or blue” combined with Rules 1 and 2 creates a classic dilemma: both possibilities lead to the same conclusion. This dilemma reasoning yields “Anne is cold.” Rule 3 then derives “Anne is rough” from cold, Rule 4 derives “Anne is young” from rough, and Rule 6 derives “Anne is nice” from young. Rule 5 creates a circular reinforcement but doesn’t alter the conclusions.
The logical structure can be represented as:
$(G_{a}\lor B_{a})\land(G_{a}\rightarrow C_{a})\land(B_{a}\rightarrow C_{a})\vdash C_{a}$
  • $G_{a}$: Green(Anne)
  • $B_{a}$: Blue(Anne)
  • $C_{a}$: Cold(Anne)
Variation 1: Rule Reduction with Same Conclusions
This variation demonstrates that removing redundant rules preserves the AI’s ability to reach the same conclusions, illustrating that fewer rules can be equally effective.
  • Facts:
    • Anne is green or blue
  • Rules:
    • Rule 1: If someone is green then they are cold. $\forall x\,(\text{Green}(x)\rightarrow\text{Cold}(x))$
    • Rule 2: If someone is blue then they are cold. $\forall x\,(\text{Blue}(x)\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
    • Rule 6: If someone is young then they are nice. $\forall x\,(\text{Young}(x)\rightarrow\text{Nice}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: T]
    • Q2: Anne is rough. True/False? [Answer: T]
    • Q3: Anne is young. True/False? [Answer: T]
    • Q4: Anne is nice. True/False? [Answer: T]
The reasoning proceeds identically to the base case: the dilemma from Rules 1-2 yields “Anne is cold,” Rule 3 yields rough, Rule 4 yields young, and Rule 6 yields nice. The removal of Rule 5 has no impact on the conclusions, demonstrating its redundancy.
Key Insight: AI systems that recognize this redundancy can simplify their reasoning processes without sacrificing accuracy, embodying the “less is more” principle.
The simplified logical structure becomes:
$(G_{a}\lor B_{a})\land(G_{a}\rightarrow C_{a})\land(B_{a}\rightarrow C_{a})\vdash C_{a}$
$C_{a}\land(C_{a}\rightarrow R_{a})\vdash R_{a}$
$R_{a}\land(R_{a}\rightarrow Y_{a})\vdash Y_{a}$
$Y_{a}\land(Y_{a}\rightarrow N_{a})\vdash N_{a}$
  • $G_{a}=\mathrm{Green}(Anne)$
  • $B_{a}=\mathrm{Blue}(Anne)$
  • $C_{a}=\mathrm{Cold}(Anne)$
  • $R_{a}=\mathrm{Rough}(Anne)$
  • $Y_{a}=\mathrm{Young}(Anne)$
  • $N_{a}=\mathrm{Nice}(Anne)$
Variation 2: Rule Equivalence with Different Conclusions
This variation replaces multiple rules with logically equivalent fewer rules, but interestingly leads to different conclusions due to the modified rule interactions.
  • Facts:
    • Anne is green or blue
  • Rules:
    • Rule A: If someone is green or blue then they are cold. $\forall x\,((\text{Green}(x)\lor\text{Blue}(x))\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: T]
    • Q2: Anne is rough. True/False? [Answer: T]
    • Q3: Anne is young. True/False? [Answer: T]
    • Q4: Anne is nice. True/False? [Answer: F]
Rule A directly captures the dilemma of Rules 1 and 2, leading to “Anne is cold” with equivalent logical force. Rules 3 and 4 then proceed as before. However, without Rule 6, we cannot derive “Anne is nice,” resulting in a different conclusion for Q4.
Key Insight: While Rule A is logically equivalent to the combination of Rules 1 and 2, the overall rule set simplification changes the available inference paths, demonstrating that equivalence at the micro-level doesn’t guarantee identical macro-level conclusions.
The logical equivalence can be shown as:
$(\forall x\,(G_{x}\rightarrow C_{x})\land\forall x\,(B_{x}\rightarrow C_{x}))\equiv\forall x\,((G_{x}\lor B_{x})\rightarrow C_{x})$
  • $G_{x}=\mathrm{Green}(x)$
  • $B_{x}=\mathrm{Blue}(x)$
  • $C_{x}=\mathrm{Cold}(x)$
However, the missing Rule 6 prevents the derivation of Nice(Anne), showing that local equivalence doesn’t preserve global derivability.
Variation 3: Rule Interference with Contradictory Conclusions
This variation adds distracting and potentially contradictory rules, testing the AI’s ability to resolve conflicts and maintain coherent reasoning.
  • Facts:
    • Anne is green or blue
    • Anne is not cold or not nice
  • Rules:
    • Rule 1: If someone is green then they are cold. $\forall x\,(\text{Green}(x)\rightarrow\text{Cold}(x))$
    • Rule 2: If someone is blue then they are cold. $\forall x\,(\text{Blue}(x)\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
    • Rule 5: If someone is young then they are cold. $\forall x\,(\text{Young}(x)\rightarrow\text{Cold}(x))$
    • Rule 6: If someone is young then they are nice. $\forall x\,(\text{Young}(x)\rightarrow\text{Nice}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: F]
    • Q2: Anne is rough. True/False? [Answer: F]
    • Q3: Anne is young. True/False? [Answer: F]
    • Q4: Anne is nice. True/False? [Answer: F]
Reasoning Process and Analysis
This variation introduces contradictory interference, which challenges the AI’s ability to resolve conflicts within the logical structure. The fact “Anne is green or blue,” combined with the original chain of reasoning (Rules 1-6), leads to the conclusions that Anne is cold, rough, young, and nice. However, the fact “Anne is not cold or not nice” introduces a conflict, as it implies that being nice would mean Anne is not cold. This creates an inconsistency that different AI systems might resolve differently:
  • Conservative approach: Detect inconsistency and withhold conclusions
  • Priority-based approach: Apply rule priorities or specificity heuristics
  • Paraconsistent approach: Accept some contradictions and continue reasoning
In this case, a conservative reasoning system would recognize the contradiction and potentially reject all derived conclusions, resulting in false for all questions.
The addition of interfering rules not only tests the AI’s ability to ignore distractions but also its capacity for inconsistency detection and resolution.
The contradiction can be formally represented as:
$(C_{a}\land N_{a})\land(N_{a}\rightarrow\neg C_{a})\vdash\bot$
  • $C_{a}=\mathrm{Cold}(Anne)$
  • $N_{a}=\mathrm{Nice}(Anne)$
  • $\bot$ represents a value that is always false.
Comparative Analysis
Table 2 provides a clear comparison of how different modifications to the rule set lead to distinct conclusion patterns, highlighting the sensitivity of reasoning systems to structural changes. These variations demonstrate how even minor adjustments to the logical framework can significantly impact the reasoning process and the final outcomes.
Variations Cold Rough Young Nice
Base Example T T T T
Rule Reduction T T T T
Rule Equivalence T T T F
Rule Interference F F F F
Table 2: Comparison of conclusions across variations

### v2

@@
 Abstract
-Large language models (LLMs) excel across many natural language tasks, yet their generalisation to structural perturbations in logical contexts remains poorly understood.
-We introduce a controlled evaluation framework that probes reasoning reliability through four targeted stress tests: (1) rule deletion, removing either redundant or essential rules from a multi-step inference chain; (2) contradictory evidence injection; (3) logic-preserving rewrites generated through several families of equivalence laws (contrapositive, double negation, implication, De Morgan, identity, and commutativity); and (4) multi-law equivalence stacking that introduces 2–5 simultaneous logical transformations.
-Across three representative model families—BERT, Qwen2, and LLaMA-like models—our experiments reveal a strikingly consistent pattern: all models achieve perfect accuracy on the base tasks and remain fully generalise to redundant rule deletion and all equivalence-based rewrites (single or multi-law), but fail sharply under essential rule deletion (dropping to 25% accuracy) and collapse completely in the presence of explicit contradictions (0% accuracy).
-These results demonstrate that LLMs possess stable invariance to semantic-preserving logical transformations, yet remain fundamentally brittle to missing or conflicting evidence.
-Our framework provides a clean diagnostic tool for isolating such reasoning failure modes and highlights persistent gaps in the logical generalisation abilities of current LLMs.
+Large language models (LLMs) achieve strong performance on many natural language tasks, yet their generalisation under structured perturbations of logical rule systems remains insufficiently characterised.
+We present a controlled evaluation framework that probes reasoning reliability through four stress tests: (1) rule deletion, removing redundant versus essential rules from a multi-step inference chain; (2) contradictory evidence injection; (3) logic-preserving rewrites based on equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity); and (4) multi-law equivalence stacking that composes 2–5 transformations.
+Across three representative model families—BERT, Qwen2, and LLaMA-like models—all models attain Acc=1.0000=1.0000 on the base split and show no degradation under redundant rule deletion.
+In contrast, essential rule deletion yields a pronounced decrease to near-chance performance, and injecting explicit contradictions reduces accuracy to 0.0000.
+Under logic-preserving rewrites, accuracy is largely preserved for single-law transformations with only small degradations in a few cases, whereas multi-law stacking exposes model-dependent sensitivity: BERT matches the base condition, TinyLlama shows only marginal degradation, and Qwen2 exhibits a substantial drop.
+Overall, the results indicate that contemporary LLMs are generally stable under semantic-preserving reformulations, yet remain brittle to missing or inconsistent evidence and may degrade under composed logical transformations depending on the model family.
+The proposed framework provides a concise diagnostic tool for isolating these failure modes and for evaluating logical generalisation beyond surface-form variation.
 Codehttps://github.com/14H034160212/lemo
 Introduction
-Reasoning over structured rule systems is a fundamental capability required for reliable decision-making and language understanding.
-Although large language models (LLMs) achieve impressive performance across numerous natural language benchmarks, it is still unclear to what extent these models perform true logical reasoning rather than relying on memorized patterns or superficial correlations.
-Much prior work evaluates LLMs in settings where all relevant information is present in canonical form; however, real-world reasoning often involves incomplete information, rephrased or transformed rules, redundant evidence, and even explicit contradictions.
-Understanding how LLMs behave under such perturbations is essential for both theoretical insight and safe deployment.
-To address this gap, we introduce a highly controlled evaluation framework that isolates core aspects of rule-based reasoning.
-The framework is built around a canonical multi-step logical chain (e.g., green → cold → rough → young → nice) that requires models to integrate several rules in sequence.
-From this base structure, we generate a family of targeted variants that probe different dimensions of reasoning generalisation:
+Reasoning over structured rule systems is a foundational capability for reliable decision-making and language understanding.
+Although large language models (LLMs) perform strongly on many benchmarks, it remains unclear to what extent such performance reflects robust logical inference as opposed to distributional heuristics.
+Moreover, practical reasoning settings frequently involve incomplete information, redundant premises, rephrased or logically transformed rules, and even explicit contradictions.
+Characterising model behaviour under these controlled perturbations is therefore important for both scientific understanding and safe deployment.
+To address this gap, we introduce a controlled evaluation framework that isolates key aspects of rule-based reasoning.
+The benchmark is built around a canonical multi-step inference chain (e.g., green $\rightarrow$ cold $\rightarrow$ rough $\rightarrow$ young $\rightarrow$ nice) requiring multi-hop integration of rules.
+From each base instance, we generate targeted variants designed to probe distinct dimensions of generalisation:
 Rule deletion.
-We remove either redundant rules—whose absence should not change conclusions—or essential rules that constitute the backbone of the inference chain.
-This tests whether models distinguish between logically important and unimportant components.
+We remove either redundant rules (whose removal should preserve derivability) or essential rules required to complete the inference chain, thereby testing sensitivity to missing inferential links.
 Contradictory evidence injection.
-We introduce explicit facts that conflict with conclusions implied by the rule chain, requiring the model to reconcile or prioritize between mutually inconsistent pieces of information.
+We inject explicit facts that conflict with rule-derived conclusions, testing whether models can accommodate inconsistency rather than defaulting to a single inference path.
 Logic-preserving transformations.
-We rewrite rules using several major families of logical equivalence laws, including contraposition, double negation, implication-to-disjunction, De Morgan’s laws, identity, and commutativity of disjunction.
-These variants preserve the underlying semantics while changing surface form, allowing us to test whether LLMs track logical content rather than specific linguistic templates.
+We rewrite rules using standard equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity).
+These variants preserve semantics while altering surface form, allowing us to test whether models track logical content beyond templated phrasing.
 Multi-law equivalence stacking.
-We introduce conditions in which 2–5 logical equivalence laws are applied simultaneously to create highly redundant, cluttered, and structurally complex rule sets, all of which remain logically equivalent to the base rules.
-We evaluate this framework across three distinct model families—BERT, Qwen2, and LLaMA-like models—using identical training and testing pipelines.
-Despite architectural differences, all models exhibit the same qualitative and quantitative behavior:
-  • Perfect reasoning on the base tasks (100% accuracy)
-  • Full generalisation to redundant rule deletion (100% accuracy)
-  • Catastrophic failure under essential rule deletion (25% accuracy)
-  • Complete collapse under explicit contradictions (0% accuracy)
-  • Full invariance to all six classes of logical equivalence rewrites (100% accuracy)
-  • Full invariance even under multi-law equivalence stacking (100% accuracy)
-Together, these results reveal a consistent pattern: LLMs correctly identify which rules are structurally necessary, generalise flawlessly across a wide spectrum of logically equivalent transformations, but lack the capacity to adjust reasoning when evidence is incomplete or contradictory.
-Rather than flexibly integrating conflicting information, models tend to complete the rule chain whenever possible, even when doing so contradicts stated facts.
-Our contributions provide a clean and extensible diagnostic tool for probing reasoning behavior, enabling precise identification of reasoning failure modes.
-The findings highlight a persistent gap between surface-level generalisation to paraphrase-like variation and broader generalisation to structural perturbation or conflict, underscoring the need for models and training strategies that can reason reliably under minimal, missing, or inconsistent evidence.
+We compose 2–5 equivalence laws simultaneously to create structurally more complex but logically equivalent rule sets, testing robustness under composed transformations.
+We evaluate the framework on three model families—BERT, Qwen2, and LLaMA-like models—under identical training and evaluation pipelines.
+Despite architectural differences, the observed behaviour is broadly consistent:
+  • Base reasoning is solved at Acc=1.0000=1.0000.
+  • Redundant rule deletion yields no measurable degradation.
+  • Essential rule deletion produces a marked decrease to near-chance performance.
+  • Injecting explicit contradictions reduces accuracy to 0.0000.
+  • Single-law equivalence rewrites are handled robustly, with only minor degradations in a small number of cases.
+  • Multi-law stacking exposes model-dependent sensitivity: some models retain baseline accuracy, whereas others exhibit substantial degradation.
+Together, these results highlight an asymmetry: contemporary LLMs are generally stable under semantic-preserving rewrites of rules, yet highly sensitive to missing or inconsistent evidence, and may degrade under composed logical transformations.
+The proposed framework provides an extensible, fine-grained diagnostic suite for isolating these failure modes, complementing existing reasoning benchmarks that often conflate multiple sources of variation.
 Related Work
@@
 Split Acc $\Delta$ Acc $\Delta$ Acc $\Delta$
-base 1.00 0.00 1.00 0.00 1.00 0.00
-variant1 1.00 0.00 1.00 0.00 1.00 0.00
-variant2 0.25 -0.75 0.25 -0.75 0.25 -0.75
-variant3 0.00 -1.00 0.00 -1.00 0.00 -1.00
-variant4-contrapositive 1.00 0.00 1.00 0.00 1.00 0.00
-variant4-double-negation 1.00 0.00 1.00 0.00 1.00 0.00
-variant4-implication 1.00 0.00 1.00 0.00 1.00 0.00
-variant4-de-morgan 1.00 0.00 1.00 0.00 1.00 0.00
-variant4-identity 1.00 0.00 1.00 0.00 1.00 0.00
-variant4-commutativity 1.00 0.00 1.00 0.00 1.00 0.00
-variant4-multi (2–5 laws) 1.00 0.00 1.00 0.00 1.00 0.00
+base 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant1 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant2 0.2950 -0.7050 0.2500 -0.7500 0.2500 -0.7500
+variant3 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
+variant4-contrapositive 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant4-double-negation 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant4-implication 1.0000 0.0000 0.9525 -0.0475 1.0000 0.0000
+variant4-de-morgan 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant4-identity 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant4-commutativity 0.9925 -0.0075 1.0000 0.0000 1.0000 0.0000
+variant4-multi (2–5 laws) 1.0000 0.0000 0.6450 -0.3550 0.9925 -0.0075
 Table 1: Accuracy and deviation from base ($\Delta$) for all models across structural variants.
-Table 1 presents a unified comparison of BERT, Qwen2, and TinyLlama across all structural variants in our benchmark.
-Despite architectural differences, all three models exhibit an almost perfectly aligned pattern of behavior.
-Each model achieves 100% accuracy on the base reasoning task and retains this performance under redundant rule removal (Variant 1) as well as all logic-preserving equivalence transformations, including six single-law rewrites and multi-law (2–5) equivalence stacking.
-These conditions yield $\Delta=0$, confirming complete invariance to semantic-preserving logical reformulations.
-In contrast, performance collapses when essential rule information is removed (Variant 2), with accuracy dropping uniformly to 25% for all models, corresponding to a $\Delta$ of 0.75-0.75.
-Even more strikingly, all models fail entirely under contradictory facts (Variant 3), scoring 0% accuracy and yielding the largest degradation from baseline ($\Delta=-1.00$).
-This cross-model consistency indicates that generalisation to logical equivalence and brittleness to structural disruptions are not tied to model size or architecture, but instead reflect systematic properties of current LLM reasoning mechanisms.
-The variations in our framework reveal key properties of LLM-based reasoning.
-By systematically manipulating the logical environment—removing rules, introducing contradictions, and applying logic-preserving equivalence transformations—we expose dimensions of generalisation and brittleness often hidden in more complex benchmarks.
-Together, these findings clarify what contemporary LLMs can and cannot generalise in rule-based reasoning.
+Table 1 reports accuracy (Acc) and deviation from the base condition ($\Delta=\mathrm{Acc}_{\mathrm{variant}}-\mathrm{Acc}_{\mathrm{base}}$) for BERT, Qwen2, and TinyLlama across all structural variants.
+All models achieve Acc=1.0000=1.0000 on the base split and exhibit no degradation under redundant rule removal (Variant 1; $\Delta=0$).
+By contrast, removing an essential rule (Variant 2) yields a substantial drop (BERT: 0.2950, $\Delta=-0.7050$; Qwen2/TinyLlama: 0.2500, $\Delta=-0.7500$), indicating strong sensitivity to missing inferential links.
+Injecting explicit contradictory facts (Variant 3) reduces accuracy to 0.0000 for all models ($\Delta=-1.0000$), suggesting that the models do not reliably revise conclusions in the presence of inconsistency.
+For logic-preserving single-law rewrites (Variant 4), performance is largely preserved: most conditions remain at 1.0000 accuracy, with only small degradations in a few settings (e.g., BERT under commutativity: 0.9925, $\Delta=-0.0075$; Qwen2 under implication-to-disjunction: 0.9525, $\Delta=-0.0475$).
+Under multi-law equivalence stacking (2–5 laws), robustness becomes model-dependent: BERT matches the base condition (1.0000), TinyLlama shows only a marginal reduction (0.9925, $\Delta=-0.0075$), whereas Qwen2 exhibits a pronounced decrease (0.6450, $\Delta=-0.3550$).
+This contrast suggests that compositional logical rewrites can introduce effective difficulty beyond single transformations, even when semantic equivalence is preserved.
 Generalisation and Sensitivity
-The progression from identical conclusions (Variation 1) to partially disrupted conclusions (Variation 2) to complete breakdown (Variation 3) reveals the fine-grained sensitivity of LLM reasoning to structural properties of the rule set.
-In Variation 1, the removal of a redundant rule leaves all models completely unaffected—indicating that LLMs can identify when a rule does not contribute new inferential content.
-In sharp contrast, Variation 2 demonstrates that eliminating a single essential step in the reasoning chain causes accuracy to collapse to chance levels, suggesting that LLMs cannot reconstruct missing inferential links even when the rest of the context is intact.
-Variation 3 highlights a different form of fragility: all models overwhelmingly prefer rule-driven inferences over explicitly stated contradictory facts.
-This behavior implies that LLMs tend to follow the internal logic of the given rule system rather than adjusting their conclusions in response to new or conflicting information—an important consideration for applications involving dynamic or noisy environments.
-The invariance observed across all logic-preserving rewrites—whether single-law or multi-law—shows that LLMs possess semantic stability, treating contraposition, double negation, implication re-expression, De Morgan transformations, and other equivalence laws as interchangeable.
-This indicates that they capture structural relationships between logically equivalent formulations, even though they falter with incomplete or inconsistent evidence.
-Overall, LLMs remain generalise to semantic-preserving transformations but extremely brittle to structural disruptions, a duality with important implications for AI safety: systems that appear reliable under surface-level variation may still fail when essential information is missing or contradictory, highlighting the need for methods that compensate for incomplete reasoning or reconcile conflicting evidence.
+Variants 1–3 isolate model sensitivity to perturbations that preserve versus disrupt derivability.
+In Variant 1 (redundant rule deletion), all models match the base accuracy, consistent with the expectation that removing unused rules should not affect entailment.
+In Variant 2 (essential rule deletion), accuracy drops to near-chance levels, suggesting limited premise tracking: models often behave as if the missing link were still present and tend to “complete” the original inference pattern even when intermediate support is removed.
+This behaviour is consistent with reliance on learned chain templates rather than verifying support for each queried conclusion.
+In Variant 3 (contradictory evidence), all models score 0.0000, indicating that they do not reliably adopt contradiction-aware conservative reasoning under the benchmark’s scoring semantics.
+For logic-preserving rewrites (Variant 4), performance is largely preserved for single-law transformations, with only minor degradations in a few cases (e.g., implication-to-disjunction or commutativity).
+However, multi-law stacking exposes model-dependent robustness: BERT remains at the base level, TinyLlama shows only marginal reduction, whereas Qwen2 degrades substantially.
+This suggests that composing multiple equivalence transformations introduces an effective distribution shift beyond isolated rewrites, likely through increased operator density and structural complexity despite semantic equivalence.
+Implications and Diagnostic Value
+The variant-wise profile provides a compact diagnostic of three failure modes: (i) missing-link sensitivity under essential rule deletion, indicating weak verification of minimal supporting premises; (ii) inconsistency handling under contradiction injection, where models fail to revise or withhold conclusions when counter-evidence is present; and (iii) compositional sensitivity, where multi-law rewrites can degrade performance even when each rewrite is semantics-preserving.
+The strong gap between single-law and multi-law conditions cautions against equating robustness to isolated perturbations with robustness under composition.
+These diagnostics motivate targeted interventions such as deletion-aware training, explicit proof/rationale verification, and stress-testing robustness under composed transformations rather than single perturbations alone.
 Conclusion
-This work introduced a controlled framework for evaluating LLM generalisation under systematic perturbations to logical reasoning.
-By constructing multi-step inference chains and generating variants involving rule deletion, contradiction injection, and multiple families of logic-preserving transformations, we isolated reasoning behaviors often obscured in prior benchmarks.
-Across BERT, Qwen2, and LLaMA-like models, we observed perfect performance and complete invariance to semantic-preserving rewrites, but sharp failures when essential rules were removed and total collapse under explicit contradictions.
-These results highlight a clear gap between surface-level generalisation and deeper logical competence, underscoring the need for models that can integrate incomplete information, handle conflicting evidence, and recover missing reasoning steps.
-Our framework provides a foundation for improving both LLM reasoning reliability and evaluation methodology.
+This work introduced a controlled framework for evaluating LLM generalisation under systematic perturbations of rule-based reasoning.
+By constructing multi-step inference chains and generating variants involving rule deletion, contradiction injection, and multiple families of logic-preserving transformations, we isolate behaviours that are often obscured in more heterogeneous benchmarks.
+Across BERT, Qwen2, and LLaMA-like models, we observe Acc=1.0000=1.0000 on the base split and no degradation under redundant rule deletion, but pronounced failures under essential rule deletion and uniform breakdown under explicit contradictions.
+For equivalence-based rewrites, performance is largely preserved for single-law transformations, while composed (multi-law) rewrites reveal model-dependent sensitivity.
+These findings underscore the gap between robustness to semantic-preserving reformulations and reliable reasoning under missing, inconsistent, or compositionally transformed evidence, and motivate methods that explicitly address premise tracking and inconsistency management.
 Appendix A Appendix
@@
     • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
-    • Rule 4: If someone is not young then they are not rough. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
+    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
     • Rule 5: If someone is young then they are cold. $\forall x\,(\text{Young}(x)\rightarrow\text{Cold}(x))$

### v3

@@
-Less Is More for Multi-Step Logical Reasoning of LLM Generalisation Under Rule Removal, Paraphrasing, and Compression
+Conflict-Aware Fusion: Resolving Logic Inertia in Large Language Models via Structured Cognitive Priors
 Abstract
-Large language models (LLMs) achieve strong performance on many natural language tasks, yet their generalisation under structured perturbations of logical rule systems remains insufficiently characterised.
-We present a controlled evaluation framework that probes reasoning reliability through four stress tests: (1) rule deletion, removing redundant versus essential rules from a multi-step inference chain; (2) contradictory evidence injection; (3) logic-preserving rewrites based on equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity); and (4) multi-law equivalence stacking that composes 2–5 transformations.
-Across three representative model families—BERT, Qwen2, and LLaMA-like models—all models attain Acc=1.0000=1.0000 on the base split and show no degradation under redundant rule deletion.
-In contrast, essential rule deletion yields a pronounced decrease to near-chance performance, and injecting explicit contradictions reduces accuracy to 0.0000.
-Under logic-preserving rewrites, accuracy is largely preserved for single-law transformations with only small degradations in a few cases, whereas multi-law stacking exposes model-dependent sensitivity: BERT matches the base condition, TinyLlama shows only marginal degradation, and Qwen2 exhibits a substantial drop.
-Overall, the results indicate that contemporary LLMs are generally stable under semantic-preserving reformulations, yet remain brittle to missing or inconsistent evidence and may degrade under composed logical transformations depending on the model family.
-The proposed framework provides a concise diagnostic tool for isolating these failure modes and for evaluating logical generalisation beyond surface-form variation.
-Codehttps://github.com/14H034160212/lemo
-Introduction
+Large language models (LLMs) excel at many natural language tasks, yet their reasoning reliability under structured perturbations of rule-based systems remains brittle.
+We present a controlled evaluation framework consisting of four stress tests: (1) rule deletion (redundant vs.
+essential); (2) contradictory evidence injection; (3) logic-preserving rewrites; and (4) multi-law equivalence stacking.
+While representative model families (BERT, Qwen2, and TinyLlama) achieve Acc=1.0000=1.0000 on base tasks, our framework reveals a critical failure mode termed Logic Inertia—a total breakdown (Acc=0.0000=0.0000) under contradictions, where deductive momentum overrides factual reality.
+To resolve this, we propose Conflict-Aware Fusion, a framework grounded in the Cognitive Structure Hypothesis which posits that robust reasoning requires an explicit structural inductive bias.
+By imposing a dual-process architecture that separates premise verification from logical deduction, Conflict-Aware Fusion successfully eliminates logic inertia, achieving 1.0000 accuracy on both base and contradictory stress tests, and significantly enhancing robustness to missing evidence.
+Our results demonstrate that, for reliable multi-step reasoning, structural verification discipline is as critical as training data scale, providing a novel blueprint for building robust, contradiction-aware AI systems  [^1] https://github.com/14H034160212/lemo .
+See the OpenAI/Evals pull request  [^2] https://github.com/openai/evals/pull/1622 .
+1 Introduction
 Reasoning over structured rule systems is a foundational capability for reliable decision-making and language understanding.
-Although large language models (LLMs) perform strongly on many benchmarks, it remains unclear to what extent such performance reflects robust logical inference as opposed to distributional heuristics.
-Moreover, practical reasoning settings frequently involve incomplete information, redundant premises, rephrased or logically transformed rules, and even explicit contradictions.
-Characterising model behaviour under these controlled perturbations is therefore important for both scientific understanding and safe deployment.
-To address this gap, we introduce a controlled evaluation framework that isolates key aspects of rule-based reasoning.
-The benchmark is built around a canonical multi-step inference chain (e.g., green $\rightarrow$ cold $\rightarrow$ rough $\rightarrow$ young $\rightarrow$ nice) requiring multi-hop integration of rules.
-From each base instance, we generate targeted variants designed to probe distinct dimensions of generalisation:
-Rule deletion.
-We remove either redundant rules (whose removal should preserve derivability) or essential rules required to complete the inference chain, thereby testing sensitivity to missing inferential links.
-Contradictory evidence injection.
-We inject explicit facts that conflict with rule-derived conclusions, testing whether models can accommodate inconsistency rather than defaulting to a single inference path.
-Logic-preserving transformations.
-We rewrite rules using standard equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity).
-These variants preserve semantics while altering surface form, allowing us to test whether models track logical content beyond templated phrasing.
-Multi-law equivalence stacking.
-We compose 2–5 equivalence laws simultaneously to create structurally more complex but logically equivalent rule sets, testing robustness under composed transformations.
-We evaluate the framework on three model families—BERT, Qwen2, and LLaMA-like models—under identical training and evaluation pipelines.
-Despite architectural differences, the observed behaviour is broadly consistent:
-  • Base reasoning is solved at Acc=1.0000=1.0000.
-  • Redundant rule deletion yields no measurable degradation.
-  • Essential rule deletion produces a marked decrease to near-chance performance.
-  • Injecting explicit contradictions reduces accuracy to 0.0000.
-  • Single-law equivalence rewrites are handled robustly, with only minor degradations in a small number of cases.
-  • Multi-law stacking exposes model-dependent sensitivity: some models retain baseline accuracy, whereas others exhibit substantial degradation.
-Together, these results highlight an asymmetry: contemporary LLMs are generally stable under semantic-preserving rewrites of rules, yet highly sensitive to missing or inconsistent evidence, and may degrade under composed logical transformations.
-The proposed framework provides an extensible, fine-grained diagnostic suite for isolating these failure modes, complementing existing reasoning benchmarks that often conflate multiple sources of variation.
-Related Work
-Research on logical reasoning in large language models (LLMs) spans multi-step deduction, logic-driven augmentation, abductive explanation, abstract reasoning assessment, generalisation analysis, and hybrid neuro-symbolic methods.
-A growing body of work shows that LLMs often fail even on elementary logical transformations.
-For example, the Reversal Curse demonstrates that models trained on “A is B” frequently fail to infer the logically equivalent “B is A,” revealing structural weaknesses in bidirectional generalisation (berglund2023reversal).
-Such findings motivate deeper investigation into how LLMs internalise logical relationships and under which conditions reasoning fidelity deteriorates.
-Early studies like Transformers as Soft Reasoners demonstrated that transformers can perform multi-step inference over natural language rules and generalise to deeper reasoning depths (transformerClark).
-Subsequent work showed that this generalisation is fragile: Bao et al.
-found that models trained on short reasoning chains fail to extrapolate to longer or structurally shifted chains, highlighting persistent deficiencies in compositional generalisation (bao2022multi).
-Similar concerns are echoed in broader analyses showing that LLMs lack strong inductive biases and struggle with systematic, relational, or abstract reasoning (ijcai2024p693; Cheng2025Em; cheng2026sapp).
-A parallel line of work attempts to strengthen reasoning via logic-driven augmentation.
+Although large language models (LLMs) perform strongly on many benchmarks, a fundamental question remains unresolved: do these models genuinely perform logical inference, or do they primarily rely on pattern completion learned from data distributions?
+This distinction is not merely theoretical.
+In real-world deployments—such as legal analysis, scientific discovery, and automated decision systems—reasoning agents must operate under conditions of incomplete information, redundant evidence, and explicit contradictions.
+A system that cannot properly validate premises before drawing conclusions risks producing confident yet logically invalid outputs.
+Consequently, understanding and improving the robustness of LLM reasoning under such conditions is both a scientific priority and a prerequisite for safe deployment.
+However, existing evaluations of LLM reasoning largely conflate linguistic competence with logical robustness, providing limited insight into how models behave when the structural integrity of rule systems is deliberately perturbed.
+In particular, the field lacks diagnostic frameworks capable of isolating the effects of missing rules, redundant rules, and contradictory premises within a unified experimental setting.
+Without such controlled analysis, it remains difficult to determine whether current models truly reason over structured knowledge or merely emulate reasoning-like behavior.
+To address this gap, we introduce a controlled evaluation framework that systematically probes the structural robustness of rule-based reasoning through four complementary stress tests: (1) Rule deletion, which distinguishes the impact of removing redundant versus essential rules; (2) Contradictory rule injection, which evaluates whether models can detect and appropriately respond to logical inconsistency; (3) Logic-preserving transformations, which assess invariance under equivalence-based reformulations of rules; and (4) Multi-law stacking, which tests compositional robustness under increasingly complex rule interactions.
+By isolating these factors within a unified benchmark, our framework provides a fine-grained diagnostic lens on the structural reliability of LLM reasoning.
+Through a comprehensive evaluation of three representative model families—BERT, Qwen2, and TinyLlama—we uncover a previously unrecognized failure pattern that we term the Asymmetry of Robustness.
+Contemporary LLMs exhibit strong stability under semantics-preserving rewrites, suggesting sensitivity to surface-level logical equivalence.
+Yet when confronted with explicit contradictions, they display a systematic collapse in performance, yielding an accuracy of 0.0000.
+We characterize this phenomenon as Logic Inertia: the tendency of models to persist along learned deductive trajectories even when foundational premises become inconsistent.
+This finding exposes a critical blind spot in current training paradigms, which appear to optimize for completing inferential chains rather than verifying their validity.
+Importantly, the universality of this failure is further confirmed through external evaluation on the Human Last Exam platform [^3] https://lastexam.ai/, where all top-tier models failed our constructed cases.
+These results suggest that the inability to resolve logical conflict is not a marginal weakness but a structural limitation of current LLM reasoning.
+Building on this diagnosis, we propose Conflict-Aware Fusion, a reasoning framework grounded in the Cognitive Structure Hypothesis.
+We argue that robust reasoning requires an explicit structural separation between premise validation and deductive execution—an inductive bias largely absent from current end-to-end training regimes.
+Our framework operationalizes this principle through a dual-process architecture in which a System 2–style contradiction detection stage precedes a System 1–style rule application stage.
+This design enforces structural verification before inference, thereby preventing deductive processes from proceeding over inconsistent premises.
+Empirical results demonstrate that this structural intervention not only resolves logic inertia but also establishes a new level of robustness across all evaluated stress conditions.
+This work makes three primary contributions.
+First, it establishes a diagnostic benchmark that enables precise measurement of structural robustness in rule-based reasoning, disentangling the effects of redundancy, incompleteness, and contradiction.
+Second, it identifies and formalizes the phenomenon of Logic Inertia, revealing a fundamental limitation in current LLM reasoning that remains invisible under conventional benchmarks.
+Third, it introduces the Conflict-Aware Fusion framework, demonstrating that incorporating explicit premise-validation mechanisms can transform model behavior, achieving 1.0000 accuracy under both standard and contradictory conditions while substantially improving resilience to essential rule deletion (Variant 2).
+More broadly, our findings challenge the prevailing assumption that scaling data and model size alone will yield reliable reasoning.
+Instead, they suggest that the next stage of progress in artificial reasoning systems depends on integrating structural verification mechanisms that enforce logical consistency prior to inference.
+By providing both a diagnostic lens and a constructive solution, this work contributes toward the development of reasoning systems that are not only powerful, but also verifiably reliable.
+2 Related Work
+Fragility of Logical Reasoning in LLMs
+Research on logical reasoning in Large Language Models (LLMs) encompasses a broad spectrum, including multi-step deduction, abductive explanation, and abstract reasoning assessment Berglund et al.
+(2023); Gendron et al.
+(2024); Cheng et al.
+(2025); Xiong et al.
+(2025); Cheng et al.
+(2026).
+While early studies suggested that transformers could act as “soft reasoners” capable of multi-step inference Clark et al.
+(2021), subsequent work has revealed significant structural weaknesses.
+For instance, the “Reversal Curse” refers to a failure of bidirectional generalization: when a model learns facts in the form “A is B,” it often cannot infer the logically equivalent reverse fact “B is A.” For example, even though “A is the mother of B” logically implies “B is the child (or son/daughter) of A,” models trained on only one direction struggle to make this inference in the reverse direction.
+Berglund et al.
+(2023).
+Furthermore, Bao et al.
+Bao et al.
+(2022) demonstrate that LLMs struggle with compositional generalization, often failing to extrapolate beyond the reasoning depths or structural patterns encountered during training.
+Logic-Driven Augmentation and Neuro-Symbolic Methods
+To address these deficiencies, a parallel line of research explores logic-driven data augmentation and hybrid architectures.
 Wang et al.
-propose a symbolic context-extension and augmentation framework that applies equivalence laws to derive new training instances, improving model performance on ReClor and LogiQA (wang-etal-2022-logic).
+(2022) proposed a symbolic context-extension framework utilizing equivalence laws to augment training sets, while Bao et al.
 Bao et al.
-extend this idea by using Abstract Meaning Representation (AMR) to generate semantically grounded logical variants that enhance generalisation to unseen reasoning depths (bao2024).
-Despite their effectiveness, these methods primarily expand data distributions rather than isolating the minimal perturbations that cause reasoning failures.
-Beyond deduction, abductive reasoning has been explored through frameworks such as AbductionRules, which train models to generate rules explaining surprising inputs (young-etal-2022-abductionrules).
-Other studies examine LLM generalisation under structured perturbations: Bao et al.
-show that reordering, compressing, or restructuring premises leads to substantial degradation in reasoning fidelity (baoAccess).
-Hybrid neuro-symbolic approaches, such as ChatLogic, integrate LLMs with Prolog-style inference mechanisms to address these limitations, although they rely heavily on external symbolic scaffolds (10650138).
-Complementing these evaluations, Bao’s dissertation provides a unified perspective on out-of-distribution generalisation, structural generalisation, and equivalence invariance by analysing architecture-agnostic failure modes across premise deletion, contradiction insertion, and logic-preserving rewrites (bao2025thesis).
-Across these threads, prior work demonstrates meaningful progress but still lacks frameworks capable of systematically disentangling generalisation to redundant versus essential evidence, contradiction sensitivity, and invariance to different classes of logical equivalence.
-These unresolved limitations motivate our controlled, fine-grained evaluation framework that isolates minimal structural perturbations and reveals predictable failure modes in multi-step reasoning across LLM architectures.
-Methodology
-1. Synthetic Logical Environment
-Base facts + rule chain generation
-$\Downarrow$
-2. Structural Variants
-V1: Redundant rule removal
-V2: Essential rule removal V3: Contradictory facts V4: Logical equivalence laws (1–5)
-$\Downarrow$
-3. Model Fine-Tuning & Inference
-BERT / Qwen2 / TinyLlama (LoRA)
-$\Downarrow$
-4. Evaluation & Analysis
-Accuracy, $\Delta_{\text{base}}$, variant breakdown
-Figure 1: Compact system architecture of our logical reasoning generalisation framework.
-Each base example is transformed into structural variations (rule deletion, contradiction injection, and logic-equivalence rewrites).
-Models are trained on base data only and evaluated across all variants to measure generalisation and sensitivity.
-The system architecture in Figure 1 outlines the full pipeline used to construct, transform, and evaluate logical reasoning tasks.
-The process begins with the generation of synthetic logical environments containing base facts and a multi-step rule chain designed to support deterministic deductive inferences.
-From each base environment, we systematically derive multiple structural variants that probe different dimensions of reasoning generalisation: (1) removal of redundant rules, (2) removal of essential rules, (3) injection of explicit contradictions, and (4) application of a wide range of logical-equivalence transformations.
-These transformed instances, together with the untouched base examples, form a comprehensive evaluation suite.
-Models such as BERT, Qwen2, and TinyLlama are fine-tuned solely on the base cases using LoRA-based parameter-efficient adaptation.
-They are then tested on all structural variants to assess reasoning generalisation under controlled perturbations.
-The evaluation module computes accuracy and deviation from the base condition ($\Delta_{\text{base}}$), enabling fine-grained analysis of model generalisation, sensitivity to missing or conflicting information, and stability under logic-preserving rewrites.
-Extended Logic Rules for Multi-Step Reasoning
-This chapter explores the reasoning capabilities of AI through a comprehensive case study that demonstrates the principle of “less is more” in logical deduction.
-We begin with a complex base example incorporating dilemma reasoning patterns, then systematically modify it in three distinct ways to reveal different aspects of AI reasoning:
-  1. 1.
-    Rule Reduction: Removing redundant rules while maintaining the same conclusions
-  2. 2.
-    Rule Equivalence: Replace multiple rules with fewer logically equivalent rules, including the laws of contraposition, implication, double negation, De Morgan, identity, and commutativity.
-  3. 3.
-    Rule Interference: Adding contradictory rules that disrupt the AI’s reasoning process.
-All examples use consistent predicates and individuals (Anne, green, blue, cold, rough, young) to facilitate comparison across variations. The example for each variation can be found in the appendix.
-Experiment
-Dataset Construction To evaluate reasoning generalisation under controlled structural perturbations, we construct a synthetic dataset consisting of multi-step logical inference chains and systematically generated variants.
-Each base example includes a fact specifying an entity’s attribute (e.g., “Anne is green or blue”), a rule chain of five inference steps (e.g., green → cold → rough → young → nice), four questions querying intermediate and final states, and binary ground-truth answers.
-From each base example, we generate eleven diagnostic variants grouped into four categories:
-Rule Deletion.
-Variant 1: removal of a logically redundant rule.
-Variant 2: removal of a logically essential rule.
-Contradictory Evidence.
-Variant 3: addition of an explicit contradiction in the facts.
-Logic-Preserving Equivalence Laws.
-Six single-law variants using distinct logical equivalence transformations: contrapositive, double negation, implication, De Morgan, identity, and commutativity.
-Multi-Law Equivalence Stacking.
-A final variant consisting of 2–5 equivalence-law transformations simultaneously applied to the base rule.
-The dataset contains 100 base groups, of which 80 are used for training and 20 for testing; all variants are applied to all base groups, resulting in a comprehensive evaluation suite with consistent ground truth.
-All transformations retain natural-language surface forms to ensure that evaluation probes reasoning rather than symbolic manipulation skills.
-Models Evaluated We evaluate three representative model families commonly used in reasoning research: BERT-base (devlin2018bert): a bidirectional encoder-only architecture.
-Qwen2-1.5B (yang2025qwen3): a decoder-only transformer representative of modern instruction-tuned LLMs.
-TinyLlama-1.1B (zhang2024tinyllama): an efficient LLaMA-like decoder model trained on open datasets.
-Training Procedure All models are fine-tuned using LoRA (Low-Rank Adaptation) to ensure parameter-efficient training and allow direct comparison across architectures.
-Training is conducted on the 80 base examples only—none of the variants are seen during training.
-Each question is treated as an independent binary classification instance.
-Input Formatting Each training instance concatenates: [facts] + [rules] + [question] with the binary label corresponding to the ground-truth answer (“T” or “F”).
-Hyperparameters To maintain comparability across models, all fine-tuning uses: Batch size: 4, Learning rate: $2\times 10^{-5}$, Epochs: 10, AdamW optimizer, Maximum sequence length: 512, and LoRA rank: 8.
-For decoder-only models (Qwen2 and TinyLlama), the tokenizer’s pad_token is set to the eos_token to avoid padding-related errors.
-Evaluation Protocol We evaluate each model on all eleven test splits: Base test set (20 unseen base examples), Rule Deletion Variants, Contradictory Evidence Variant, Six Equivalence-Law Variants, and Multi-Law Variant.
-Each test contains four questions per example, producing 80 questions for the base set and 400 questions per variant. For each input, the model predicts a binary label.
-During evaluation, we additionally record: The exact facts and rules presented, the equivalence laws used (if any), Question-level predictions, and a human-readable description of the structural perturbation.
-This allows granular error analysis of different reasoning stressors.
-We use accuracy as the primary metric, reflecting the percentage of correctly answered questions in each test split.
-Because each question is binary and independent, accuracy provides a direct measure of reasoning correctness.
-Discussion and Implications
+(2024) leveraged Abstract Meaning Representation (AMR) to generate semantically grounded logical variants.
+Beyond data-centric approaches, neuro-symbolic methods like ChatLogic, Wang et al.
+(2024) integrate LLMs with external Prolog-style engines to ensure deductive validity.
+However, these solutions primarily focus on expanding seen distributions or relying on external symbolic scaffolds, rather than resolving the model’s intrinsic sensitivity to minimal structural changes in the input.
+Generalisation under Structured Perturbations
+Recent investigations have pivoted toward analyzing model behavior under controlled perturbations, such as reordering, compressing, or restructuring premises Bao et al.
+(2025); Young et al.
+(2022); Xiong et al.
+(2026).
+These studies indicate that reasoning fidelity degrades sharply when the surface form of a problem is altered, even if the underlying logic remains invariant.
+Building on this, Bao’s unified perspective on out-of-distribution generalization Bao (2025) identifies architecture-agnostic failure modes across premise deletion, contradiction injection, and logical rewrites.
+Despite these advances, there remains a critical need for a framework capable of systematically disentangling sensitivity to redundant versus essential evidence, contradiction handling, and invariance across diverse equivalence laws.
+This gap motivates our fine-grained evaluation framework, which isolates minimal structural perturbations to reveal predictable failure patterns in multi-step reasoning.
+3 Methodology
+We introduce the Conflict-Aware Fusion framework, a structurally guided reasoning approach that enforces an explicit separation between premise verification and deductive execution.
+Guided by the Cognitive Structure Hypothesis, our methodology integrates three interdependent components: (1) a structural robustness benchmark for controlled evaluation, (2) a dual-process reasoning architecture that enforces verification-before-deduction, and (3) a structure-aligned optimization pipeline.
+Together, these components operationalize reasoning as a structurally constrained process rather than a purely data-driven behavior.
+3.1 Logic Inertia as a Design Constraint
+As established in Section 1, contemporary LLMs exhibit logic inertia: a tendency to persist along learned deductive trajectories even when foundational premises are incomplete or contradictory.
+In this work, logic inertia is treated not merely as an empirical observation but as an operational design constraint.
+Existing approaches that rely on contradiction exposure or large-scale data augmentation do not enforce an explicit verification stage within the reasoning trajectory, allowing deductive continuation to proceed unchecked.
+We therefore formalize the Cognitive Structure Hypothesis:
+Reliable multi-step reasoning requires an explicit structural separation between premise verification and deductive execution.
+Under this hypothesis, robust reasoning depends on the structural organization of inference rather than solely on knowledge scale or exposure.
+Models must first establish the logical coherence and sufficiency of premises before applying inference rules.
+This principle directly informs both the evaluation framework and the architecture introduced below.
+3.2 Structural Robustness Benchmark
+A central component of our methodology is a structural robustness benchmark designed for controlled and fine-grained evaluation of reasoning reliability.
+The benchmark serves a dual role: it functions as a diagnostic tool for identifying structural reasoning failures and as a structured data source for model training and optimization.
+Unlike conventional benchmarks composed of heterogeneous tasks, all evaluation instances are derived from a single canonical rule-based system with systematically controlled perturbations.
+Each variant preserves domain semantics and vocabulary while differing only in structural properties, enabling behavioral differences to be attributed directly to reasoning robustness rather than domain shift.
+Canonical reasoning backbone.
+The canonical backbone defines a reference reasoning state from which all perturbed variants are derived.
+A concrete instantiation of this backbone is provided in the Appendix as the base example, which illustrates specific facts, rules, and deduction chains.
+This distinction clarifies that the benchmark is not limited to this single instance, but systematically evaluates reasoning over a unified structural template.
+Controlled structural perturbations.
+Benchmark splits are generated by modifying one structural property at a time while preserving semantic domain and vocabulary. Structural robustness is evaluated along four orthogonal dimensions:
+  • Structural necessity sensitivity: preserve conclusions under redundant rule deletion while revising conclusions when essential rules are removed.
+  • Consistency verification: detect logical contradictions prior to inference.
+  • Logical-form invariance: maintain identical entailments under semantics-preserving transformations.
+  • Compositional robustness: remain stable under stacked structural rewrites.
+Benchmark generation (Variant 2 & 3).
+We generate all instances from a single canonical rule-based template (canonical backbone) and create hard structural variants by controlled perturbations that preserve surface vocabulary and domain semantics.
+Given a base instance $(F,R,Q)$ (facts, rules, and a fixed set of yes/no queries), we construct: Variant 2 by removing a key rule on the main deduction chain, which breaks downstream entailments; and Variant 3 by injecting an explicit contradictory fact, which forces the model to detect inconsistency and halt/reject deduction.
+All splits are serialized into a unified CSV schema containing facts, rules, questions, and answers.
+Other variants are deferred to the Appendix.
+Algorithm 1 Structural Variant Generation (Variant 2 & 3)
+1:Canonical generator $\mathsf{Base}(\cdot)$, key rule $r^{\star}$, contradiction template $\mathsf{Contr}(\cdot)$, number of instances NN
+2:Datasets $\mathcal{D}_{base}$, $\mathcal{D}_{v2}$, $\mathcal{D}_{v3}$
+3:for $i=1$ to NN do
+4:  $(F,R,Q,A)\leftarrow\mathsf{Base}(i)$ $\triangleright$ canonical backbone instance
+5:  $\mathcal{D}_{base}\leftarrow\mathcal{D}_{base}\cup\{(F,R,Q,A)\}$
+6:  $R_{v2}\leftarrow R\setminus\{r^{\star}\}$ $\triangleright$ Variant 2: remove an essential rule
+7:  $A_{v2}\leftarrow\mathsf{Label}(F,R_{v2},Q)$
+8:  $\mathcal{D}_{v2}\leftarrow\mathcal{D}_{v2}\cup\{(F,R_{v2},Q,A_{v2})\}$
+9:  $F_{v3}\leftarrow F\cup\{\mathsf{Contr}(F,R)\}$ $\triangleright$ Variant 3: inject a contradiction
+10:  $A_{v3}\leftarrow\mathsf{Label}(F_{v3},R,Q)$
+11:  $\mathcal{D}_{v3}\leftarrow\mathcal{D}_{v3}\cup\{(F_{v3},R,Q,A_{v3})\}$
+12:end for
+13:Serialize $\mathcal{D}_{base},\mathcal{D}_{v2},\mathcal{D}_{v3}$ to CSV with a unified schema.
+3.3 Architectural Overview: Dual-Process Reasoning
+Guided by the structural failures identified above, we introduce the Conflict-Aware Fusion architecture.
+The core innovation is the imposition of a dual-process reasoning structure within the Chain-of-Thought (CoT) generation path.
+As illustrated in Figure 1, the reasoning process is constrained such that deductive rules cannot be applied until the model has explicitly validated the logical consistency of the premises.
+This design converts verification from an optional behavior into a required structural step.
+When contradictions or missing dependencies are detected, the model halts or revises reasoning rather than continuing forward deduction.
+Conflict-Aware CoT Input Facts: Socrates is a man.
+Rule: All men are mortal.
+New Fact: Socrates is NOT mortal.
+Instruction: ”Step 1: Check Contradiction…” SFT Logic Verify?
+Trace A (Correct): “Step 1: Conflict Verified!
+(Man $\to$ Mortal vs Not Mortal) Step 2: STOP.
+Trace B (Naive): “Step 1: Consistent.
+Step 2: Deduce Mortal.” Result: False (Halt) Result: True (Hallucinated) $\pi_{DPO}$ Verification $\pi_{SFT}$ HallucinationSystem 2 CheckSystem 1 Inertia
+Figure 1: Conflict-Aware Fusion Architecture.
+The Prompt Template (top) imposes a structural prior requiring an explicit “Contradiction Detection” (Step 1).
+The model generates reasoning traces that branch based on verification.
+DPO (left box) reinforces the “Halt” path when contradictions are present, resolving Logic Inertia.
+This architecture operationalizes the Cognitive Structure Hypothesis by embedding a verification-first inductive bias directly into the reasoning process, ensuring that inference is based on validated premises.
+3.4 Structure-Aligned Optimization Pipeline
+To ensure consistent adherence to the proposed structural prior, we implement a two-stage optimization pipeline that explicitly reinforces verification-before-deduction as a learned reasoning constraint.
+Stage 1: Structural SFT.
+We first perform supervised fine-tuning on a balanced dataset of 11,200 instances spanning canonical, perturbed, and contradiction-containing variants.
+All training samples enforce a mandatory “Step 1: Verify facts” preamble, requiring the model to explicitly assess premise completeness and consistency before attempting deduction.
+By uniformly applying this structure across both base and robustness-focused variants, the training process discourages shortcut learning on familiar patterns and promotes generalizable verification behavior.
+This stage establishes premise checking as a default procedural step rather than an optional response to rare anomalies.
+Stage 2: Logical Alignment via DPO.
+We further refine the model using Direct Preference Optimization (DPO) to align generation behavior with structural reliability.
+Preference pairs are constructed to contrast verification-compliant reasoning with opportunistic deductive continuation.
+Responses that correctly trigger a “Halt Reasoning” state upon detecting contradictions or insufficient premises are consistently preferred over those that proceed with unsupported inference.
+This alignment step directly penalizes hallucinated shortcuts and reinforces disciplined termination when logical validity cannot be established.
+Through this combined pipeline, premise verification is internalized as a governing structural constraint that shapes how reasoning unfolds rather than merely influencing final answers.
+The model learns not only to produce correct conclusions but to regulate the conditions under which deduction is permitted, thereby reducing logic inertia and improving robustness under structural perturbations.
+4 Experiment and Results
+4.1 Main Results: Conflict-Aware Fusion Performance
+The primary objective of our evaluation is to assess whether the proposed framework preserves logical integrity under controlled structural stressors.
+As reported in Table 1, all baseline models indicate strong performance under unperturbed conditions.
+However, this apparent reliability collapses once the structural integrity of the rule system is disrupted.
+When essential rules are removed (Variant 2), accuracy drops sharply across all models, and under explicit contradictions (Variant 3), performance uniformly falls to 0.0000.
+These results provide direct evidence of pronounced logic inertia.
 BERT Qwen2 TinyLlama
@@
 variant1 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
-variant2 0.2950 -0.7050 0.2500 -0.7500 0.2500 -0.7500
-variant3 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
-variant4-contrapositive 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
-variant4-double-negation 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
-variant4-implication 1.0000 0.0000 0.9525 -0.0475 1.0000 0.0000
-variant4-de-morgan 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
-variant4-identity 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
-variant4-commutativity 0.9925 -0.0075 1.0000 0.0000 1.0000 0.0000
-variant4-multi (2–5 laws) 1.0000 0.0000 0.6450 -0.3550 0.9925 -0.0075
-Table 1: Accuracy and deviation from base ($\Delta$) for all models across structural variants.
-Table 1 reports accuracy (Acc) and deviation from the base condition ($\Delta=\mathrm{Acc}_{\mathrm{variant}}-\mathrm{Acc}_{\mathrm{base}}$) for BERT, Qwen2, and TinyLlama across all structural variants.
-All models achieve Acc=1.0000=1.0000 on the base split and exhibit no degradation under redundant rule removal (Variant 1; $\Delta=0$).
-By contrast, removing an essential rule (Variant 2) yields a substantial drop (BERT: 0.2950, $\Delta=-0.7050$; Qwen2/TinyLlama: 0.2500, $\Delta=-0.7500$), indicating strong sensitivity to missing inferential links.
-Injecting explicit contradictory facts (Variant 3) reduces accuracy to 0.0000 for all models ($\Delta=-1.0000$), suggesting that the models do not reliably revise conclusions in the presence of inconsistency.
-For logic-preserving single-law rewrites (Variant 4), performance is largely preserved: most conditions remain at 1.0000 accuracy, with only small degradations in a few settings (e.g., BERT under commutativity: 0.9925, $\Delta=-0.0075$; Qwen2 under implication-to-disjunction: 0.9525, $\Delta=-0.0475$).
-Under multi-law equivalence stacking (2–5 laws), robustness becomes model-dependent: BERT matches the base condition (1.0000), TinyLlama shows only a marginal reduction (0.9925, $\Delta=-0.0075$), whereas Qwen2 exhibits a pronounced decrease (0.6450, $\Delta=-0.3550$).
-This contrast suggests that compositional logical rewrites can introduce effective difficulty beyond single transformations, even when semantic equivalence is preserved.
-Generalisation and Sensitivity
+variant2 (Essential Deletion) 0.2950 -0.7050 0.2500 -0.7500 0.2500 -0.7500
+variant3 (Contradiction) 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
+variant4-multi (Stacked Laws) 1.0000 0.0000 0.6450 -0.3550 0.9925 -0.0075
+Table 1: Baseline Performance: Accuracy and deviation ($\Delta$) across structural variants without fusion optimization.
+To resolve these failure modes, we propose the Fusion-Conflict framework.
+As summarized in Table 2, our method sets a new benchmark for logical reasoning, achieving 1.0000 on both the base task and the contradiction stress test (Variant 3).
+Table 2: Final Performance: Comparison of All Methods (Accuracy). Fusion-Conflict (Ours) achieves 100% accuracy on contradictions while recovering base performance.
+Method Base Acc Var 2 (Rule Removal) Var 3 (Contradiction) Rank
+Stage 1 (SFT Baseline) 0.512 0.250 0.210 7
+DPO (Direct Alignment) Rafailov et al. (2023) 0.475 0.267 0.510 6
+CoT (Standard) Wei et al. (2022) 0.500 0.390 0.865 5
+Mixed-Aug (Data Augmentation without CoT) 0.525 0.405 0.972 4
+RA-CoT (Standard CoT + DPO) 0.263 0.593 0.690 3
+Fusion-LRA (Conflict-Aware SFT) 0.988 0.753 0.705 2
+[…]
+4.2 Experimental Setup and Hyperparameters
+Dataset and Models The benchmark consists of multi-step inference chains with 100 base groups (80 for training, 20 for testing).
+We evaluate BERT-base, Qwen2-1.5B, and TinyLlama-1.1B.
+All training uses Low-Rank Adaptation (LoRA) for parameter efficiency.
+Stage 1 Training Details For the initial rule-prediction phase, the hyperparameters used for the generative models are detailed in Table 3.
+Hyperparameter Value
+Epochs 3
+Learning Rate $2.0\times 10^{-5}$
+Batch Size 4
+Max Sequence Length 512
+Optimizer AdamW (Mixed Precision)
+LoRA Config r: 8, $\alpha$: 16, dropout: 0.05
+Table 3: Detailed Stage 1 Training Hyperparameters.
+4.3 Ablation and Incremental Optimization
+To understand the evolution of our methodology, we track the performance of intermediate strategies.
+Table 4 highlights the impact of Stage 1 pre-training alone, which shows that while rule prediction helps slightly with Variant 2, it causes significant degradation in base task accuracy without subsequent fusion stages.
+Ablation: Effect of Stage 1 Pre-training on Generalisation
+BERT Qwen2 TinyLlama
+Split Acc $\Delta$ Acc $\Delta$ Acc $\Delta$
+base 1.0000 0.0000 0.5250 -0.4750 0.5375 -0.4625
+variant2 (Essential Deletion) 0.2500 -0.7500 0.4050 -0.1200 0.5325 -0.0050
+variant3 (Contradiction) 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
+Table 4: Ablation Results: Stage 1 pre-training moderately improves Variant 2 robustness for generative models but severely degrades base performance.
+Finally, Table 5 illustrates the trade-offs encountered during the development of Fusion-LRA.
+This intermediate step achieved an optimal balance (Rank 1 at the time) by recovering 98.8% base accuracy, providing the foundation for the final Conflict-Aware architecture.
+Method Base Acc Var 2 Var 3 Rank
+Mixed-Aug 0.525 0.405 0.972 3
+RA-CoT 0.263 0.593 0.690 2
+Fusion-LRA 0.988 0.753 0.705 1
+Table 5: Iterative Comparison: Fusion-LRA vs earlier iterations.
+4.4 Is DPO Necessary? (Impact Analysis)
+A key question in our optimization pipeline is the specific contribution of Stage 2 Reinforcement Learning (DPO) versus Supervised Fine-Tuning (SFT). Our results provide a definitive answer.
+As shown in Table 2, Fusion-LRA (SFT Only) achieved a strong Base Accuracy of 0.988 but plateaued at 0.705 for Contradiction handling (Variant 3).
+This indicates that while SFT can teach the model the form of the contradiction check, the model still occasionally “hallucinates” a successful check even when a contradiction exists (a phenomenon we term “shallow alignment”).
+In contrast, Fusion-Conflict (SFT + DPO) pushes the Variant 3 accuracy to 1.0000.
+By explicitly penalizing trajectories where the model fails to halt on contradictions (using DPO preference pairs), we sharpen the decision boundary of the Contradiction Detection.
+Consequently, DPO does not merely incrementally improve performance; it acts as the critical robustness hardener, bridging the final gap from “mostly consistent” to “logically sound.”
+5 Discussion and Implications
+5.1 Comparative Analysis of Robustness Strategies
+Our experimental results, summarized in Table 2, reveal a clear evolutionary path from baseline models to the proposed Conflict-Aware Fusion framework.
+Initially, baseline models (Stage 1) demonstrated high task proficiency but were essentially ”logic-blind” to contradictions, scoring near 0.210 on Variant 3.
+While Mixed-Aug specialized in identifying contradictions (reaching 0.972), it failed to scale to complex reasoning chains.
+The RA-CoT framework introduced more rigor but suffered from an ”over-conservatism” trap, where the model preferred to output ”False” for nearly everything, causing Base Accuracy to plummet to 0.263.
+The Fusion-Conflict approach successfully breaks this trade-off.
+By achieving 1.0000 accuracy in both Base and Variant 3 (Table 2), it proves that the model no longer relies on opportunistic deduction.
+Instead, it has internalized the ”Contradiction Detection” as a prerequisite for logic, effectively solving the ”Logic Inertia” problem that plagued all previous iterations.
+5.2 The Success of the Cognitive Structure Hypothesis
+The decisive performance of Conflict-Aware Fusion provides strong empirical support for our Cognitive Structure Hypothesis.
+As demonstrated by the comparison between Fusion-LRA and Fusion-Conflict, simply improving data balance (LRA) leads to substantial performance gains (0.705 on V3); however, these gains remain inferior to those achieved by explicitly modeling cognitive conflict.
+This suggests that for LLMs, ”Reasoning” is not just a statistical continuation of tokens, but a process that can be steered by structural inductive biases.
+The ”Step 1: Verify facts” preamble acts as a functional circuit breaker, preventing the model’s System 1 (intuition/pattern matching) from overriding the factual reality of the premises.
+5.3 Generalisation and Sensitivity
 Variants 1–3 isolate model sensitivity to perturbations that preserve versus disrupt derivability.
-In Variant 1 (redundant rule deletion), all models match the base accuracy, consistent with the expectation that removing unused rules should not affect entailment.
-In Variant 2 (essential rule deletion), accuracy drops to near-chance levels, suggesting limited premise tracking: models often behave as if the missing link were still present and tend to “complete” the original inference pattern even when intermediate support is removed.
-This behaviour is consistent with reliance on learned chain templates rather than verifying support for each queried conclusion.
-In Variant 3 (contradictory evidence), all models score 0.0000, indicating that they do not reliably adopt contradiction-aware conservative reasoning under the benchmark’s scoring semantics.
-For logic-preserving rewrites (Variant 4), performance is largely preserved for single-law transformations, with only minor degradations in a few cases (e.g., implication-to-disjunction or commutativity).
-However, multi-law stacking exposes model-dependent robustness: BERT remains at the base level, TinyLlama shows only marginal reduction, whereas Qwen2 degrades substantially.
-This suggests that composing multiple equivalence transformations introduces an effective distribution shift beyond isolated rewrites, likely through increased operator density and structural complexity despite semantic equivalence.
-Implications and Diagnostic Value
-The variant-wise profile provides a compact diagnostic of three failure modes: (i) missing-link sensitivity under essential rule deletion, indicating weak verification of minimal supporting premises; (ii) inconsistency handling under contradiction injection, where models fail to revise or withhold conclusions when counter-evidence is present; and (iii) compositional sensitivity, where multi-law rewrites can degrade performance even when each rewrite is semantics-preserving.
-The strong gap between single-law and multi-law conditions cautions against equating robustness to isolated perturbations with robustness under composition.
-These diagnostics motivate targeted interventions such as deletion-aware training, explicit proof/rationale verification, and stress-testing robustness under composed transformations rather than single perturbations alone.
-Conclusion
-This work introduced a controlled framework for evaluating LLM generalisation under systematic perturbations of rule-based reasoning.
-By constructing multi-step inference chains and generating variants involving rule deletion, contradiction injection, and multiple families of logic-preserving transformations, we isolate behaviours that are often obscured in more heterogeneous benchmarks.
-Across BERT, Qwen2, and LLaMA-like models, we observe Acc=1.0000=1.0000 on the base split and no degradation under redundant rule deletion, but pronounced failures under essential rule deletion and uniform breakdown under explicit contradictions.
-For equivalence-based rewrites, performance is largely preserved for single-law transformations, while composed (multi-law) rewrites reveal model-dependent sensitivity.
-These findings underscore the gap between robustness to semantic-preserving reformulations and reliable reasoning under missing, inconsistent, or compositionally transformed evidence, and motivate methods that explicitly address premise tracking and inconsistency management.
+In Variant 1 (redundant rule deletion), our Fusion model maintains Acc=1.000, confirming that the internal contradiction detection does not introduce fragility to noise.
+However, Variant 2 (Essential Rule Removal) remains the most challenging stress test.
+While Fusion-Conflict achieves a high score of 0.735, it still shows a delta compared to the base set.
+This indicates that while the model is excellent at detecting explicit contradictions (V3), it is still slightly prone to ”logical leaps” when an inferential link is missing but the overall pattern looks familiar.
+This diagnostic highlights a future research direction: enhancing the model’s sensitivity to logical gaps as much as it is now sensitive to logical conflicts.
+5.4 Implications for Reliable AI Deployment
+The transition from RA-CoT’s over-conservatism to Fusion-Conflict’s precision has significant implications for high-stakes deployment:
+  1. 1.
+    Process over Data: Improving reasoning reliability is better achieved by refining the reasoning process (via structured CoT) than by merely expanding variables in the training set.
+  2. 2.
+    DPO as a Logical Refiner: Our use of DPO (Stage 2) demonstrates that preference alignment can be used to punish hallucinatory logic, shifting the model from ”answering” to ”verifying.”
+  3. 3.
+    Diagnostic Value: Our multi-variant benchmark allows researchers to distinguish between a model that ”knows the rules” and a model that ”knows when the rules don’t apply.”
+5.5 Real-World Validation: LogicNLI & MNLI
+To verify the generalization of our Cognitive Structure Hypothesis, we extended our evaluation to real-world datasets, specifically LogicNLI Tian et al.
+(2021) (a First-Order Logic NLI benchmark) and the contradiction subset of MNLI Williams et al.
+(2018).
+Unlike synthetic rules, these datasets feature diverse linguistic patterns and implicit logical structures.
+We fine-tuned a lightweight 0.5B model (Qwen2.5-0.5B) using our Conflict-Aware Template (Stage 1 SFT).
+The model was taught to explicitly output “Step 1: Detect Contradiction” before predicting entailment or contradiction.
+Initial results are highly promising: the model successfully learned to invoke the “Halt” mechanism upon detecting conflicts in LogicNLI samples.
+Method LogicNLI MNLI (Con)
+Standard SFT (Baseline) 58.5% 34.1%
+Mixed-Aug (Data Only) 65.2% 41.5%
+Fusion-LRA (No DPO) 92.4% 78.3%
+[…]
+Table 6: Zero-shot transfer performance.
+“MNLI (Con)” refers to accuracy specifically on the contradiction subset of the Multi-Genre NLI dataset, which tests the model’s ability to reject inconsistent premises in open-domain text.
+This demonstrates that the specific structural prior we propose—separating verification from deduction—is not limited to synthetic logic but is a viable strategy for robust natural language inference in complex domains.
+6 Conclusion
+In this work, we present a controlled framework for analyzing and enhancing the structural robustness of rule-based reasoning in large language models.
+By applying systematically designed perturbations, we demonstrate that contemporary models—including BERT, Qwen2, and TinyLlama—maintain high accuracy under standard and form-equivalent conditions, yet consistently fail when essential rules are removed or explicit contradictions are introduced.
+This behavior reflects a persistent reliance on forward inferential continuation rather than explicit premise validation.
+To address this limitation, we introduce the Conflict-Aware Fusion framework, which enforces an explicit separation between premise verification and deductive execution within the reasoning trajectory.
+Empirical results show that this structural intervention substantially improves robustness under rule deletion and fully resolves contradiction-induced failure, achieving 1.0000 accuracy in contradiction handling while preserving base-task performance.
+These findings provide direct empirical support for the Cognitive Structure Hypothesis: reliable multi-step reasoning depends not only on data exposure but also on the structural organization of the reasoning process.
+Beyond these technical results, our work demonstrates that the gap between semantic-form invariance and logical-content reliability can be systematically bridged through architectural priors in Chain-of-Thought prompts.
+By operationalizing explicit premise verification, Conflict-Aware Fusion offers a scalable blueprint for doubt-aware AI systems capable of disciplined, evidence-driven inference.
+Overall, this work contributes both a diagnostic benchmark and a constructive methodology for developing LLMs that reason with verifiable logical rigor rather than surface-level pattern completion.
 Appendix A Appendix
-Base Example: Complex Dilemma Reasoning Structure
+A.1 Base Example: Complex Dilemma Reasoning Structure
 We begin with a comprehensive example that establishes the core reasoning pattern, drawing on a structure analogous to the Paradox of the Court [^4] The Paradox of the Court involves a contract between the teacher Protagoras and his student Euathlus, where the student only pays for lessons if he wins a court case.
@@
   • $C_{a}$: Cold(Anne)
-Variation 1: Rule Reduction with Same Conclusions
+A.2 Variation 1: Rule Reduction with Same Conclusions
 This variation demonstrates that removing redundant rules preserves the AI’s ability to reach the same conclusions, illustrating that fewer rules can be equally effective.
@@
   • $N_{a}=\mathrm{Nice}(Anne)$
-Variation 2: Rule Equivalence with Different Conclusions
+A.3 Variation 2: Rule Equivalence with Different Conclusions
 This variation replaces multiple rules with logically equivalent fewer rules, but interestingly leads to different conclusions due to the modified rule interactions.
@@
 However, the missing Rule 6 prevents the derivation of Nice(Anne), showing that local equivalence doesn’t preserve global derivability.
-Variation 3: Rule Interference with Contradictory Conclusions
+A.4 Variation 3: Rule Interference with Contradictory Conclusions
 This variation adds distracting and potentially contradictory rules, testing the AI’s ability to resolve conflicts and maintain coherent reasoning.
@@
     • Q4: Anne is nice. True/False? [Answer: F]
-Reasoning Process and Analysis
+A.5 Reasoning Process and Analysis
 This variation introduces contradictory interference, which challenges the AI’s ability to resolve conflicts within the logical structure.
@@
 However, the fact “Anne is not cold or not nice” introduces a conflict, as it implies that being nice would mean Anne is not cold.
-This creates an inconsistency that different AI systems might resolve differently:
-  • Conservative approach: Detect inconsistency and withhold conclusions
+This creates an contradiction that different AI systems might resolve differently:
+  • Conservative approach: Detect contradiction and withhold conclusions
   • Priority-based approach: Apply rule priorities or specificity heuristics
@@
 In this case, a conservative reasoning system would recognize the contradiction and potentially reject all derived conclusions, resulting in false for all questions.
-The addition of interfering rules not only tests the AI’s ability to ignore distractions but also its capacity for inconsistency detection and resolution.
+The addition of interfering rules not only tests the AI’s ability to ignore distractions but also its capacity for contradiction detection and resolution.
 The contradiction can be formally represented as:
@@
   • $\bot$ represents a value that is always false.
-Comparative Analysis
+A.6 Comparative Analysis
 Table 7 provides a clear comparison of how different modifications to the rule set lead to distinct conclusion patterns, highlighting the sensitivity of reasoning systems to structural changes.
@@
 Table 7: Comparison of conclusions across variations
+Split BERT Qwen2 TinyLlama
+Acc $\Delta$ Acc $\Delta$ Acc $\Delta$
+base 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant1 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant2 0.2950 -0.7050 0.2500 -0.7500 0.2500 -0.7500
+variant3 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
+variant4-contrapositive 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant4-double-negation 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant4-implication 1.0000 0.0000 0.9525 -0.0475 1.0000 0.0000
+variant4-de-morgan 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant4-identity 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
+variant4-commutativity 0.9925 -0.0075 1.0000 0.0000 1.0000 0.0000
+variant4-multi (2–5 laws) 1.0000 0.0000 0.6450 -0.3550 0.9925 -0.0075
+Table 8: Accuracy and deviation from base ($\Delta$) for all models across structural variants.
+Table 8 reports accuracy (Acc) and deviation from the base condition ($\Delta=\mathrm{Acc}_{\text{variant}}-\mathrm{Acc}_{\text{base}}$) for BERT, Qwen2, and TinyLlama across all structural variants.
+All models achieve Acc = 1.0000 on the base split and exhibit no degradation under redundant rule removal (Variant 1; $\Delta=0$).
+By contrast, removing an essential rule (Variant 2) yields a substantial drop (BERT: 0.2950, $\Delta=-0.7050$; Qwen2/TinyLlama: 0.2500, $\Delta=-0.7500$), indicating strong sensitivity to missing inferential links.
+Injecting explicit contradictory facts (Variant 3) reduces accuracy to 0.0000 for all models ($\Delta=-1.0000$), suggesting that the models do not reliably revise conclusions in the presence of inconsistency.
+Figure 2: An example where all top-tier models failed the Human Last Exam.
+Figure 2 shows a screenshot of the AGI Safe AI dashboard under the section “II.
+AI Answers”, where multiple state-of-the-art language models (including Claude Sonnet 4.5, GPT-4.1, GPT-5.2, Claude Opus 4.5, and Gemini 3 Pro Preview) are evaluated on a logical reasoning question involving a set of premises about “Anne” and four sub-questions (Q1–Q4).
+This example 2 evaluates reasoning with a disjunctive color premise and a conflicting constraint.
+From the fact “Anne is green or blue” and the rules stating that being green or being blue implies being cold, it follows that Anne is cold in either case.
+Applying the rule “if someone is cold then they are rough,” we can further derive that Anne is rough.
+However, the additional fact “Anne is not cold or not nice” introduces a conflict with the derived conclusion that Anne is cold.
+Given $\mathrm{Cold}(\mathrm{Anne})$, the disjunction $\neg\mathrm{Cold}(\mathrm{Anne})\lor\neg\mathrm{Nice}(\mathrm{Anne})$ forces the conclusion that Anne is not nice.
+As a result, the statement that Anne is nice is contradicted.
+Regarding youth, the rules only support forward implications: being young implies being cold and being nice.
+There is no rule that allows inferring that Anne is young from coldness, roughness, or any other established fact.
+Therefore, Anne being young cannot be derived from the given information.
+Different reasoning frameworks may handle the inconsistency differently (e.g., conservative systems may block some inferences, whereas paraconsistent systems may continue reasoning in the presence of contradictions).
+Under standard conservative reasoning, Anne is cold and rough are supported, Anne is young is undetermined, and Anne is nice is contradicted.
+The core interaction can be summarized as:
+$C_{a}\land(\neg C_{a}\lor\neg N_{a})\vdash\neg N_{a}$
+  • $C_{a}=\mathrm{Cold}(\mathrm{Anne})$
+  • $N_{a}=\mathrm{Nice}(\mathrm{Anne})$
+Each model’s output is displayed in separate panels with its predicted answers and reasoning.
+The results indicate that the models produce inconsistent and conflicting conclusions, with some models claiming all statements are true while others detect inconsistency in the premises.
+At the bottom, the evaluation summary reports that 0 out of 5 models answered correctly, suggesting that this example is challenging even for top-tier models and highlights their limitations in formal logical reasoning and consistency checking.